OpenStudy (anonymous):

Given the sequence: {2, 9, 16, 23, 30…}, a30 = ???

OpenStudy (anonymous):

what should i use to solve this

OpenStudy (anonymous):

firstly find the recursive formula

OpenStudy (anonymous):

let \(a_0=2\) \(a_1=9\) \(a_2=16\) \(a_3=23\) ....

OpenStudy (anonymous):

retriceu can see \[16=a_2=2*9-2=2a_1-a_0\]\[a_3=2a_2-a_1\]...\[a_n=?\]

OpenStudy (anonymous):

7

OpenStudy (anonymous):

7 ? what?

OpenStudy (anonymous):

a

OpenStudy (anonymous):

so a30=212

OpenStudy (anonymous):

is it right?

OpenStudy (anonymous):

yes it is... \[a_n=2a_{n-1}-a_{n-2}\] characterstic equation\[s^2-2s+1=0\] \(s=1\) with multiplicity of 2 so we can write \(a_n=A+Bn\) find A and B u will get A=2 , B=7 \(a_n=2+7n\) so \(a_{30}=2+7*30=212\)